OpenStudy (anonymous):

2-6i/3+i perform the indicated operation and simplify, write the result in form a+bi????

OpenStudy (anonymous):

just multiply by conjugate

OpenStudy (anonymous):

(2-6i)/(3+i) * (3-i)/(3-i)

OpenStudy (lgbasallote):

\[\frac{2-6i}{3+i}\] multiply by the njugate \[\frac{2-6i}{3+i} \times \frac{3-i}{3-i}\] \[\frac{(2-6i)(3-i)}{(3+i)(3-i)} \implies \frac{(2-6i)(3+i)}{9 - i^2} \implies \frac{(2-6i)(3+i)}{9-(-1)} \implies \frac{(2-6i)(3+i)}{10}\] now all that's left is simplify (2-6i)(3+i) and you'll get the answer

OpenStudy (lgbasallote):

do you still need help?

OpenStudy (anonymous):